# General Preference Modeling with Preference Representations for Aligning Language Models

### v1

@@ 5.2 General Preference Optimization (GPO)
$\max_{\bm{\theta}}\ \mathbb{E}_{\mathbf{x}}\left[\mathbb{E}_{\mathbf{y}\sim\pi_{\bm{\theta}}(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\mu(\cdot\mid\mathbf{x})}\left[s(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x})\right]\right]-\beta\mathbb{E}_{\mathbf{x}}\left[\mathrm{KL}\left(\pi_{\bm{\theta}}(\cdot\mid\mathbf{x})\|\pi_{\text{ref}}(\cdot\mid\mathbf{x})\right)\right],$ (5.5)
where $\pi_{\text{ref}}$ is a reference policy (e.g., the initial language model), $\mu$ is the opponent policy (usually the same as $\pi_{\text{ref}}$), and $\beta>0$ is a regularization parameter controlling the divergence from the reference policy. We would like to point out that this formulation is different from the many previous works (Wu et al., 2024b; Swamy et al., 2024; Rosset et al., 2024; Munos et al., 2023; Azar et al., 2023) as they consider maximizing the win rate $\mathbb{P}(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})$, while our formulation is to maximize $s(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})=\log\frac{\mathbb{P}(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})}{\mathbb{P}(\mathbf{y}\prec\mathbf{y}^{\prime}|\mathbf{x})}$. Note that $\mathbb{P}(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})$ only varies between 00 and $1$, while $s(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})$, similar to the reward $r(\mathbf{y};\mathbf{x})$ in RLHF or DPO, can take arbitrary values. The flexibility in its value range might benefit fine-tuning.
General Preference Optimization. We consider the SPPO loss used by Wu et al. (2024b) for iterative preference optimization, except that we use preference score instead of preference probability in the loss form. SPPO used $K$ responses for each prompt $\mathbf{x}$ and calculated the empirical win rate of each response $\mathbf{y}_{k}$. Instead, we calculate $\widehat{s}\left(\mathbf{y}_{i}\succ\mu\mid\mathbf{x}\right)$ to estimate the empirical win rate over the distribution $\mu$ as below:
@@
$\bm{\theta}_{t+1}=\arg\min_{\bm{\theta}}\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}_{t}}(\cdot\mid\mathbf{x})}\Bigg{[}\left(\log\left(\frac{\pi_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}\mid\mathbf{x})}\right)-\frac{1}{\beta}\left(\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)-\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\right)\right)^{2}\Bigg{]},$ (5.7)
where the normalizing factor $Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x}):=\sum_{\mathbf{y}}\pi_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})\exp\left(\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)\right)$.
In practice, we directly replace $\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})$ with 00111In late stages of the iterative training, $\pi_{\bm{\theta}_{t}}$ is close to equilibrium so the preference model can not distinguish between policy $\pi_{\bm{\theta}}$ and the opponent policy $\pi_{\bm{\theta}_{t}}$( meaning $\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)\approx 0$). Therefore, we have $\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\approx 0$.. Intuitively, if a response $\mathbf{y}$ receives a high average score, GPO will increase its log probability. We report the empirical performance of GPO in Section 6.3.
@@
Notice that the GPO learning objective can be seen as an offline policy gradient algorithm (see Appendix C) for the optimization problem defined in Equation (5.7), similar results have been discussed in Munos et al. (2023); Wu et al. (2024b).
Remark 5.2.
Note that the general preference score given by our GPM in Equation (5.5) can also be integrated as preference (reward) signal for any off-the-shelf RLHF and preference optimization methods, including (iterative) DPO (Rafailov et al., 2024), IPO (Azar et al., 2023), NLHF (Munos et al., 2023), SPPO (Wu et al., 2024b) and REBEL (Gao et al., 2024), as well as PPO-based methods (Ouyang et al., 2022) by directly optimizing Equation (5.5).
@@ 6 Experiments
We conducted extensive experiments to evaluate the effectiveness of the proposed General Preference representation model (GPM) in comparison to traditional reward-based models, particularly focusing on its ability to model cyclic preferences and improve language model alignment. Our experiments are designed to address the following key questions:
  • Q1: Can the GPM effectively capture and model cyclic and intransitive preferences, where traditional models like the Bradley-Terry (BT) reward model struggle?
  • Q2: How does the GPM perform on standard preference modeling benchmarks (RewardBench) compared to the BT model?
  • Q3: How does using the GPM for downstream policy optimization impact language model performance on real-world tasks compared to reward-based approaches?
6.1 Cyclic Preference Modeling
To address Q1, we evaluate the ability of the GPM to capture intransitive, cyclic preferences that traditional transitive models (like the BT model) struggle to represent.
Cyclic Preference Dataset. We constructed a dataset by inducing cyclic preferences from the Ultrafeedback dataset Cui et al. (2024). The dataset includes responses evaluated across four key metrics: instruction following, honesty, truthfulness, and helpfulness. We created preference cycles such as: instruction following $\succ$ honesty $\succ$ truthfulness $\succ$ helpfulness $\succ$ instruction following, ensuring the presence of intransitive cycles. We further generated four sub-datasets by omitting one metric from each cycle, resulting in datasets of varying complexity with 216 to 363 instances.
Training and Evaluation. We trained the GPM using the Gemma-2B-it language model as the base and evaluated the models based on their ability to predict the human-provided preferences in these datasets. Since cyclic preferences are inherently intransitive, we measure accuracy as the percentage of correctly predicted human preferences, where higher scores indicate better handling of non-transitive preferences.
Table 1: Comparison of Bradley-Terry (BT) reward model and General Preference representation models (GPM) on cyclic preference datasets.
@@ 6.1 Cyclic Preference Modeling
BT RM No instruction following 62.4
GPM No instruction following 100.0 (+37.6)
BT RM No honesty 61.6
GPM No honesty 100.0 (+38.4)
BT RM No truthfulness 50.0
GPM No truthfulness 100.0 (+50.0)
BT RM No helpfulness 62.9
GPM No helpfulness 100.0 (+37.1)
Results and Analysis. As shown in Table 1, the GP representation model achieves 100% accuracy across all datasets, significantly outperforming the BT model. These results validate the GP representation model’s ability to capture complex, cyclic preferences, confirming the theoretical advantages of using a preference representation-based approach over traditional reward models that assume transitivity.
6.2 Experiments on RewardBench
To address Q2, we compare the GP representation model and the BT reward model on the RewardBench benchmark (Lambert et al., 2024), which covers diverse preference modeling tasks, including Chat, Chat-Hard, Safety, and Reasoning.
Datasets and Experimental Setup. We train both the BT and GPMs using the Skywork Reward Data Collection (Liu & Zeng, 2024), which contains 80,000 pairwise preference examples from tasks in various domains. We evaluate both models on RewardBench, using two different base models: Gemma-2B-it (Team et al., 2024) (2B parameters) and Llama-3.1-8B-Instruct (Dubey et al., 2024) (8B parameters), which are well-suited for instruction-following tasks (see Appendices A.2 and B.2 for the implementation details and experimental setup).
Results and Analysis. Table 2 presents the results. The GPM consistently outperforms the BT model for both base models on RewardBench, with notable improvements in tasks involving complex reasoning (e.g., Chat-Hard and Reasoning). These results highlight the superior expressiveness of the GPM in preference modeling.
Table 2: Comparison between the Bradley-Terry (BT) models and the General Preference representation models (GPM) with varying embedding head dimensions on RewardBench. The highest scores are in bold and the second highest are underlined.
@@ 6.2 Experiments on RewardBench
BT RM 1 89.7 87.3 91.0 96.2 91.1
GPM 2 92.7 87.3 90.5 96.1 91.7
4 93.6 87.1 90.4 96.0 91.8
6 93.3 88.6 90.6 96.0 92.1 ( +1.0)
8 92.7 86.4 90.4 96.7 91.6
Ablation Studies.
We further conducted ablation studies to assess the impact of varying the representation (embedding) dimension in the GPM. Table 2 shows that increasing the embedding dimension generally improves performance. Additional experimental results regarding the ablation on the scale gate and L2 normalization on the embedding head can be found in Appendix B.1 (see Appendix A for implementation details on the scale gate and the embedding head).
To address Q3, we investigate the effectiveness of the GPM in language model for alignment using Self-Play Policy Optimization (SPPO) (Wu et al., 2024b) and our proposed General Preference Optimization (GPO).
Experimental Setup. We fine-tuned language models using SPPO and GPO, integrating preference scores provided by our GP representation model (GPM). We evaluated the models on AlpacaEval 2.0 (Dubois et al., 2024) and MT-Bench (Zheng et al., 2023), two widely used benchmarks for evaluating LLM alignment.
Results and Analysis. The evaluation results on the benchmarks are as follows. For AlpacaEval 2.0, we compared the generated responses of the aligned models with those of GPT-4-turbo. To avoid the preference bias when using GPT-4-turbo as the evaluator, we also used DeepSeek-V2 (DeepSeek-AI, 2024) and GPT-4o-mini as the evaluators besides GPT-4-turbo itself. Notice that the Length Controlled (LC) Win Rate results are using a generalized linear model fitted using default evaluator GPT-4-turbo, so it does not apply to other evaluators. The results of the three evaluators are presented in Tables 34 and 5.
For MT-Bench, we used the default mode to let GPT-4 grade and give a score to the model’s answer, and the MT-Bench scores of aligned models are presented in Table 6. These results suggest that integrating the General Preference representation model (GPM) into policy optimization can enhance the downstream performance of language models, supporting alignment with complex human preferences (for additional experimental results on GSM8K, MMLU, etc., please see Appendix B.1).
Table 3: AlpacaEval 2.0 evaluation results. Base model: LLama3-8B-it, Evaluator: GPT-4o-mini. The results are grouped by the size and type of the RM or PM, and the number of iterations.
Size Type Iter SPPO GPO
Win Rate Avg. Len Win Rate Avg. Len
base 32.26 1959 32.26 1959
2B BT RM 1 44.48 1843 49.55 1837
2 59.89 2028 57.19 2048
3 71.11 2244 67.81 2245
GPM 1 52.64 (+8.16) 2102 58.49 (+8.94) 2145
2 66.01 (+6.12) 2318 72.22 (+15.03) 2404
3 72.97 (+1.86) 2490 77.11 (+9.30) 2613
8B BT RM 1 40.96 1802 42.56 1764
2 52.43 1945 57.19 2080
3 55.30 1832 62.79 3445
GPM 1 45.80 (+4.84) 1878 49.55 (+6.99) 1877
2 56.04 (+3.61) 2020 57.54 (+0.35) 2126
3 60.73 (+5.43) 1994 66.23 (+3.44) 2157
Table 4: AlpacaEval 2.0 evaluation results. Base model: LLama3-8B-it, Evaluator: GPT-4-turbo. The results are grouped by the size and type of the RM or PM, and the number of iterations.
@@ 6.3 Downstream Performance on Aligning Language Models with Human Preferences
2B BT RM 1 33.71 30.88 1843 36.28 33.17 1837
2 37.38 37.88 2028 37.87 38.89 2048
3 38.86 42.98 2244 38.79 42.80 2245
GPM 1 32.64 35.12 2102 35.23 38.10 2145
2 35.16 41.40 2318 36.04 44.47 2404
3 35.30 45.44 2490 38.51 48.30 2613
8B BT RM 1 33.38 29.90 1802 35.96 31.71 1764
2 38.83 37.82 1945 41.48 41.73 2080
3 40.55 37.09 1832 42.16 43.33 3445
GPM 1 33.39 31.49 1878 37.45 35.17 1877
2 37.35 37.60 2020 39.07 39.94 2126
3 39.72 39.38 1994 41.19 43.38 2157
Table 5: AlpacaEval 2.0 evaluation results. Base model: LLama3-8B-it, Evaluator: DeepSeek-V2. The results are grouped by the size and type of the RM or PM, and the number of iterations.
Size Type Iter SPPO GPO
Win Rate Avg. Len Win Rate Avg. Len
base 30.20 1959 30.20 1959
2B BT 1 43.00 1843 43.78 1837
2 51.78 2028 52.80 2048
3 60.99 2244 58.21 2245
GP 1 48.41 (+5.41) 2102 52.57 (+8.79) 2145
2 58.12 (+6.34) 2318 62.52 (+9.72) 2404
3 65.26 (+4.27) 2490 68.28 (+10.07) 2613
8B BT 1 41.73 1802 43.36 1764
2 50.24 1945 52.90 2080
3 50.76 1832 56.59 3445
GP 1 45.42 (+3.69) 1878 46.89 (+3.53) 1877
2 52.77 (+2.53) 2020 53.03 (+0.13) 2126
3 54.73 (+3.97) 1994 59.43 (+2.84) 2157
Table 6: MT-Bench evaluation results. Base model: LLama3-8B-it, Evaluator: GPT-4.
Size Type Iter SPPO GPO
1st 2nd Avg. 1st 2nd Avg.
base 8.27 7.38 7.82 8.27 7.38 7.82
2B BT RM 1 8.11 7.44 7.77 8.15 7.37 7.77
2 8.37 7.59 7.98 8.20 7.69 7.95
3 8.26 7.36 7.81 8.59 7.54 8.06
GPM 1 8.26 7.66 7.96 8.49 7.50 7.99
2 8.26 7.66 7.96 8.38 7.55 7.97
3 8.31 7.53 7.92 8.63 7.89 8.26
8B BT RM 1 8.53 7.85 8.19 8.54 8.05 8.29
2 8.33 7.91 8.12 8.16 7.55 7.85
3 8.05 7.66 7.86 8.12 7.76 7.94
GPM 1 8.27 7.60 7.93 8.30 7.70 8.00
2 8.46 7.71 8.08 8.11 7.65 7.88
3 8.14 7.59 7.87 8.22 7.63 7.93

### v2

@@ 5 General Preference Optimization
$\max_{\bm{\theta}}\ \mathbb{E}_{\mathbf{x}}\left[\mathbb{E}_{\mathbf{y}\sim\pi_{\bm{\theta}}(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\mu(\cdot\mid\mathbf{x})}\left[s(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x})\right]\right]$
$-\beta\mathbb{E}_{\mathbf{x}}\left[\mathrm{KL}\left(\pi_{\bm{\theta}}(\cdot\mid\mathbf{x})\|\pi_{\text{ref}}(\cdot\mid\mathbf{x})\right)\right],$ (5.1)
where $\pi_{\text{ref}}$ is a reference policy (e.g., the initial language model), $\mu$ is the opponent policy (usually the same as $\pi_{\text{ref}}$), and $\beta>0$ is a regularization parameter controlling the divergence from the reference policy. We would like to point out that this formulation is different from the many previous works (Wu et al., 2024b; Swamy et al., 2024; Rosset et al., 2024; Munos et al., 2023; Azar et al., 2023) as they consider maximizing the win rate $\mathbb{P}(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})$, while our formulation is to maximize $s(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})=\log\frac{\mathbb{P}(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})}{\mathbb{P}(\mathbf{y}\prec\mathbf{y}^{\prime}|\mathbf{x})}$. Note that $\mathbb{P}(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})$ only varies between 00 and $1$, while $s(\mathbf{y}\succ\mathbf{y}^{\prime}|\mathbf{x})$, can be seen as a generalized version of the reward $r(\mathbf{y};\mathbf{x})$ in RLHF or DPO (see Section 4.3), can take arbitrary values.
General Preference Optimization (GPO). We consider the iterative preference optimization process such as SPPO (Wu et al., 2024b), while we use preference score instead of preference probability in the loss form. SPPO used $K$ responses for each prompt $\mathbf{x}$ and calculated the empirical win rate of each response $\mathbf{y}_{k}$. Instead, we calculate $\widehat{s}\left(\mathbf{y}_{i}\succ\mu\mid\mathbf{x}\right)$ to estimate the empirical win rate over the distribution $\mu$ as below:
@@
$\bm{\theta}_{t+1}=$ $\arg\min_{\bm{\theta}}\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}_{t}}(\cdot\mid\mathbf{x})}\Bigg{[}\bigg{(}\log\frac{\pi_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}\mid\mathbf{x})}-\frac{1}{\beta}\big{(}\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)-\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\big{)}\bigg{)}^{2}\Bigg{]},$ (5.3)
where we have the normalizing factor $Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x}):=\sum_{\mathbf{y}}\pi_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})\exp\left(\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)\right)$. In practice, we directly replace $\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})$ with 00111In late stages of the iterative training, $\pi_{\bm{\theta}_{t}}$ is close to equilibrium so the preference model can not distinguish between policy $\pi_{\bm{\theta}}$ and the opponent policy $\pi_{\bm{\theta}_{t}}$( meaning $\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)\approx 0$). Therefore, we have $\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\approx 0$..
Intuitively, if a response $\mathbf{y}$ receives a high average score, GPO will increase its log probability. We report the empirical performance of GPO in Section 6.3. The following theorem establishes the convergence properties of GPO:
Theorem 5.1.
Consider the optimization problem defined by the GPO loss (5.3) and assume it is realizable. Let $\{\pi_{\bm{\theta}_{t}}\}_{t=1}^{T}$ denote the sequence of policies generated by GPO, and define $\bar{\pi}_{T}=\frac{1}{T}\sum_{t=1}^{T}\pi_{\bm{\theta}_{t}}$ as the average policy. Given that the preference score $s$ is bounded within $[-\rho,\rho]$, by setting $\beta=\Theta\left(\sqrt{T}\right)$, we have:
$\max_{\pi}s\left(\pi\succ\bar{\pi}_{T}\right)-\min_{\pi}s\left(\pi\prec\bar{\pi}_{T}\right)=O\left(\frac{1}{\sqrt{T}}\right).$
Connection to Policy Gradient. Applying policy gradient on (5.1) gives:
$\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}}}\bigg{[}\bigg{(}\widehat{s}(\mathbf{y}\succ\pi_{\bm{\theta}_{t}})-\beta\log\frac{\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\bigg{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(\mathbf{y}|\mathbf{x})\bigg{]}$
So Equation (5.3) can also be seen as a policy gradient method for the optimization problem (5.1).
@@
Note that the general preference score given by our GPM can also be integrated as a preference (reward) signal for many other off-the-shelf RLHF and preference optimization methods, including (iterative) DPO-based methods (Rafailov et al., 2024), IPO (Azar et al., 2023), NLHF (Munos et al., 2023), SPPO (Wu et al., 2024b) and REBEL (Gao et al., 2024), as well as PPO-based methods (Ouyang et al., 2022) by directly optimizing problem (5.1).
@@ 6 Experiments
We conducted several experiments to evaluate the effectiveness of the proposed General Preference embedding Model (GPM) in comparison to traditional reward-based models, particularly focusing on its ability to model general preference and improve language model alignment.
Table 1: Comparison of Bradley-Terry (BT) reward models and General Preference embedding models (GPM) on cyclic preference datasets. Cyclic No. $1$: Honest $\succ$ Truthful $\succ$ Helpful $\succ$ Honesty; Cyclic No. $2$: IF $\succ$ Truthful $\succ$ Helpful $\succ$ IF; Cyclic No. $3$: IF $\succ$ Honesty $\succ$ Helpful $\succ$ IF; Cyclic No. $4$: IF $\succ$ Honesty $\succ$ Truthful $\succ$ IF.
@@
BT RM Cyclic No. $1$ 62.4
GPM Cyclic No. $1$ 100.0 (+37.6)
BT RM Cyclic No. $2$ 61.6
GPM Cyclic No. $2$ 100.0 (+38.4)
BT RM Cyclic No. $3$ 50.0
GPM Cyclic No. $3$ 100.0 (+50.0)
BT RM Cyclic No. $4$ 62.9
GPM Cyclic No. $4$ 100.0 (+37.1)
Table 2: Comparison between the Bradley-Terry (BT) models and the General Preference embedding models (GPM) with varying embedding head dimensions on RewardBench. The highest scores are in bold. Note that BT RM is a special case of GPM when embedding dimension $d=1$ (see Section 4.3).
@@
BT RM 1 88.55 85.75 91.49 96.47 90.56
GPM 2 91.62 88.38 90.68 94.82 91.37
4 93.30 86.18 91.22 95.69 91.60
6 91.90 87.50 91.62 96.40 91.86
8 93.58 87.50 91.08 95.44 91.90 ( +1.34)
Table 3: AlpacaEval 2.0 evaluation results. Base model: Llama3-8B-it, Evaluator: GPT-4-turbo. The results are grouped by the size and type of the RM or PM, and the number of iterations. Bold entries indicate that GPM outperforms BT RM under the same training settings.
@@
2B BT RM 1 31.95 31.59 1939 34.01 33.08 1929
2 36.00 36.77 2032 38.90 39.90 2049
3 40.01 42.12 2136 42.21 44.20 2151
GPM 1 30.87 32.48 (+0.89) 2066 35.27 37.95 (+4.87) 2102
2 34.54 40.76 (+3.99) 2301 36.77 42.96 (+3.06) 2343
3 36.06 45.61 (+3.49) 2498 37.74 48.25 (+4.05) 2582
8B BT RM 1 32.20 27.83 1740 36.32 30.37 1702
2 39.75 36.95 1868 41.79 40.11 1933
3 42.55 40.92 1948 40.37 38.56 1969
GPM 1 33.48 30.85 (+3.02) 1861 36.00 33.19 (+2.82) 1850
2 37.93 38.38 (+1.43) 2029 40.81 42.80 (+2.69) 2115
3 39.45 41.64 (+0.72) 2385 38.98 41.54 (+2.98) 3249
6.1 Experiments on RewardBench
We compare the GPM and BT reward model on the RewardBench benchmark (Lambert et al., 2024), which covers diverse preference modeling tasks, including Chat, Chat-Hard, Safety, and Reasoning.
Datasets and Experimental Setup. We train both BT RMs and GPMs using the decontaminated version of Skywork Reward Data Collection (Liu & Zeng, 2024), which contains around 80k pairwise preference data from tasks in various domains. We evaluate both models on RewardBench, using two different base models: Gemma-2B-it (Team et al., 2024) (2B parameters) and Llama-3.1-8B-Instruct (Dubey et al., 2024) (8B parameters), which are well-suited for instruction-following tasks (please refer to Appendix E.3 for the implementation details).
Results and Analysis. The results are presented in Table 2. On RewardBench, using the Gemma-2B-it base model, GPM achieves an average score of 82.29%, which is an improvement of 7.44% over the BT model’s average score of 74.85%. Specifically, in the Chat task, GPM improves performance from 67.32% (BT RM) to 79.61%, and in the Chat-Hard task, from 63.37% to 75.66%. For the Llama-3.1-8B-Instruct base model, GPM achieves an average score of 91.90% (embedding dimension 8), representing a 1.34% improvement over the BT model’s average score of 90.56%. In the Chat task, GPM improves from 88.55% (BT RM) to 93.58%, and in the Chat-Hard task, from 85.75% to 88.38%. Using the Gemma-2-9B-it base model, GPM achieves an average score of 92.33% (embedding dimension 4), showing an improvement of 0.87% over the BT model’s average score of 91.46%. In the Chat task, GPM boosts performance from 91.62% (BT RM) to 93.58%, and in Chat-Hard, from 85.96% to 87.72%. These results indicate that GPM outperforms the BT model across various base models and tasks, particularly in the Chat and Chat-Hard tasks. Note that BT RM is a special case of GPM when the embedding dimension $d=1$ (see Section 4.3).
Ablation studies. We conducted ablation studies to assess the impact of varying the embedding dimension in GPM. As shown in Table 2, the performance of GPM varies with the embedding dimension. For the Llama-3.1-8B-Instruct base model, an embedding dimension of 8 achieves the highest average score of 91.90%, compared to 91.86% with a dimension of 6 and 91.60% with a dimension of 4. In the Chat-Hard task with the same base model, the highest score of 88.38% is achieved with an embedding dimension of 2, compared to 87.50% with dimension 8. In addition, we can find that for the Gemma-2B-it base model, the highest average score of 82.29% is achieved with an embedding dimension of 6, showing an improvement over lower dimensions, such as 80.43% with dimension 4. These results suggest that the optimal embedding dimensions vary across different base models and tasks. For additional ablation studies on GPM architecture design, please refer to Appendix E.1.
6.2 Cyclic Preference Modeling
We evaluate the ability of GPM to capture intransitive, cyclic preferences that traditional transitive models (like the BT model) struggle to represent. Specifically, we evaluate GPMs and BT RMs on CyclicPreference datasets, which are constructed based on the Ultrafeedback dataset (Cui et al., 2024) (See Appendix E).
Training and Evaluation. We trained GPMs and BT RMs using the Gemma-2B-it language model as the base and evaluated the models based on their ability to predict intransitive preferences. For GPM, the loss function is Equation (A.1). For the Bradley-Terry (BT) model, the loss function is $\mathcal{L}=-\log\sigma(r_{w}-r_{l})$ (Ouyang et al., 2022). Since cyclic preferences are inherently intransitive, we measure accuracy as the percentage of correctly predicted human preferences, where higher scores indicate better handling of non-transitive preferences. As shown in Table 1, the GP representation model achieves near-perfect accuracy across all datasets, significantly outperforming the BT model (we report the test accuracy on the training dataset but with different comparison pairs used in the training dataset). These results validate GPM’s ability to capture complex, cyclic preferences, confirming the theoretical advantages of using a preference embedding-based approach over traditional reward models that assume transitivity.
We further investigate the effectiveness of GPM in language model for alignment using Self-Play Policy Optimization (SPPO) (Wu et al., 2024b) and our proposed General Preference Optimization (GPO), integrating preference scores provided by our GP representation model (GPM). We evaluated the models on AlpacaEval 2.0 (Dubois et al., 2024), MT-Bench (Zheng et al., 2023), GSM8K, MMLU, etc., several widely used benchmarks for evaluating LLM alignment.
Results and Analysis. The evaluation results on the benchmarks are as follows. For AlpacaEval 2.0, we compared the generated responses of the aligned models with those of GPT-4o-mini and GPT-4-turbo. The results of the three evaluators are presented in Tables 3 and 5. From Table 3, we observe that both SPPO and GPO demonstrate improved win rates with successive iterations, highlighting the iterative nature of these optimization methods, and GPO consistently outperforms SPPO. In addition, the bolded entries indicate that GPM-integrated methods consistently outperforms BT RM-based methods under the same settings on Win Rate.